OpenStudy (anonymous):

the right circular cylinder, radius of base x, height y is inscribed in a right circular cone, radius of base r and height h. express y as a function of x

OpenStudy (anonymous):



OpenStudy (anonymous):

h/y=r/x (concept of similar triangles) y=hx/r